### v1

# Beamformed 360° Sound Maps: U‑Net‑Driven Acoustic Source Segmentation and Localization

Abstract
We introduce a U-net model for 360° acoustic source localization formulated as a spherical semantic segmentation task. Rather than regressing discrete direction-of-arrival (DoA) angles, our model segments beamformed audio maps (azimuth × elevation) into regions of active sound presence. Using delay-and-sum (DAS) beamforming on a custom 24-microphone array, we generate signals aligned with drone GPS telemetry to create binary supervision masks. A modified U-Net, trained on frequency-domain representations of these maps, learns to identify spatially distributed source regions while addressing class imbalance via the Tversky loss. Because the network operates on beamformed energy maps, the approach is inherently array-independent and can adapt to different microphone configurations without retraining from scratch. The segmentation outputs are post-processed by computing centroids over activated regions, enabling robust DoA estimates. Our dataset includes real-world open-field recordings of a DJI Air 3 drone, synchronized with 360° video and flight logs across multiple dates and locations. Experimental results show that U-net generalizes across environments, providing improved angular precision, offering a new paradigm for dense spatial audio understanding beyond traditional Sound Source Localization (SSL).
Index Terms:
sound-source localization, beamforming, semantic segmentation, U-Net, drone acoustics.1 Introduction
SSL is a fundamental task in spatial audio analysis, with applications ranging from surveillance, security, search and rescue, environmental monitoring, and wildlife tracking [1]. SSL is often reduced to estimating the DoA of sound sources. The DoA is usually defined as the azimuth and elevation angle of the direction of the audio source while ignoring the distance to it. Microphone arrays can be steered to act as spatial filters, which enables manipulation of the array’s directivity (also referred to as beamforming).[2]. Traditional SSL techniques rely on signal processing algorithms such as time-difference of arrival (TDOA), generalized cross-correlation with phase transform (GCC-PHAT)[3], Steered-response power-phase transform (SRP-PHAT)[4] and multiple signal classification (MUSIC)[5], or beamforming methods like (DAS)[6]. While effective in controlled environments, these approaches degrade in performance under noise, reverberation, or when dealing with moving sources and complex acoustic scenes. Recent deep learning methods have achieved notable improvements in accuracy and robustness, even in challenging scenarios with noise, reverberation, and multiple simultaneous sources [7]. In recent years, the use of multichannel audio processing and visual perception has gained traction [8], particularly with the use of microphone arrays and beamforming to create spatial energy maps [9] essentially turning sound into images, some of those systems are known as acoustic cameras.[10].
The Acoustic ”imaging” enables the use of powerful computer vision architectures, especially convolutional neural networks (CNNs), to perform spatial reasoning on sound scenes. However, most deep learning-based SSL models outputs direction-of-arrival (DOA) angles or coordinates, typically via classification or regression. Few have explored frame-based spatial segmentation of the full acoustic field analogous to semantic segmentation in images.
In this work, we propose a novel spherical segmentation framework for localizing sound sources using microphone arrays. Inspired by the image recognition paradigm in computer vision like YOLO [11] or Deeplapv3[12], we develop a U-Net-based architecture[13] that performs binary segmentation over a 2D spherical acoustic image (azimuth × elevation) derived from DAS beamforming. Instead of regressing point estimates of direction, our model learns to segment the acoustic field, highlighting regions in space where sound sources are present. This formulation enables spatial mapping of targets such as drones. To enable supervised learning, we construct and release a labeled dataset based on real-world recordings of a DJI Air 3 drone in open-field conditions on different days and locations. The public dataset comprises 24‑channel audio, GPS‑aligned binary masks and 360° video.
Our approach provides a scalable way to learn acoustic semantic segmentation, supporting generalization to other sound sources beyond drones. Potential applications include drone detection and tracking, acoustic camera visualization, multi-source scene understanding, and real-time sound field monitoring.
The main contributions of this paper include the introduction of a novel paradigm for acoustic source localization formulated as a spherical acoustic image segmentation task, inspired by semantic segmentation approaches in computer vision. The authors design and train a U-Net-based model, which performs segmentation of sound source directions on beamformed audio maps. To support further research in audio localization and learning-based sound source localization (SSL), they also release a unique multi-channel, real-world dataset that includes synchronized audio, spatial labels, and drone telemetry.
2 Related Work
Acoustic perception tasks explicitly distinguish three sequential objectives: detection, classification, and localization. Many classical SSL systems tackle them in isolation rather than in an integrated pipeline [1]. Recent Sound Event Detection and Localisation (SELD) systems, also referred to as SSL, encompass two tasks: sound event detection (SED) and DoA, which are separate outputs of the neural network. The SED branch performs a multi-label classification task, and the DoA branch performs a multi-output regression task, as cited in [14] and [15].
Our work resolves SELD as a semantic-segmentation problem; the resulting mask simultaneously answers: (i) is there a source? (detection), (ii) where is it? (localisation) and through class specific training (iii) what type is it? (classification).
2.1 Classical Sound‑Source Localisation (SSL)
In early SSL systems, DAS beamformer was the central processing block because it is algorithmically simple, computationally light, and easy to implement in hardware [16]. Later refinements improved robustness under real-world conditions: GCC-PHAT and SRP-PHAT introduce phase-based spectral weighting to combat reverberation [17, 18], while the Minimum Variance Distortionless Response (MVDR) beamformer assigns microphone-specific weights that minimise noise without distorting the desired direction [19].
High‑resolution sub‑space methods go a step further. MUSIC [5] and ESPRIT [20] give sharper peaks, but assume narrow‑band, non‑coherent sources and perfect array calibration.
In practice, classical methods do a good job at localising a single source in low‑noise scenes, and with thresholding they can give a basic detection flag. They do not, however, offer classification. Performance also drops when the number of active sources grows or when the room is highly reverberant [21].
2.2 Deep Learning (DL) for SSL
Large audio datasets and cheaper GPUs have made deep learning attractive for SSL [1]. Different deep learning architectures have been explored:
CNNs learn spatial–spectral patterns directly from spectrograms or MFCCs [22].
RNNs and gated variants track moving sources by modelling temporal context [23].
Graph Neural Networks capture the geometry of distributed microphone arrays [24].
U‑Net family: Encoder–decoder U‑Nets have become popular because skip connections recover lost details in downsampling and give rich per-pixel output. Lee et al. reach sub‑degree accuracy for overlapping sources by correcting the problems associated with DAS beam at low frequency and suppressing side‑lobes at high frequency [27]. Building on this idea, Zhou et al. introduced audio‑visual segmentation (AVS) in which a TPAVI‑conditioned U‑Net injects audio cues at every scale to produce pixel‑wise masks of the visible sounding objects [28]. Other works add a second head so that the same network performs sound‑event detection and localisation (SELD) simultaneously, while Dense‑U‑Net further extends the concept to dynamic, high‑noise audio‑visual scenes [29]. Unlike AVS, our approach removes the dependency on vision, thereby enabling efficacy even when the source lies outside the camera’s field‑of‑view or under poor visibility.
Deep networks often produce false positives when trained exclusively on segments that contain sources. [30] show that incorporating silence frames into the training set improves robustness to background noise and prevents ghost detections. Inspired by this finding, we augment our drone dataset with “no‑drone” recordings, enabling the Tversky‑loss to direct the U‑Net to learn a calibrated decision boundary between the presence and absence of sound source of interest.
2.3 Gap Analysis and Motivation
Speed vs. Robustness. Classical DAS‑based methods offer low computational latency but lose resolution and struggle with heavy noise or many sources. Fully trained DL models can be robust to noise and can detect, localize, and classify, but they usually require large training sets.
Hybrid path. By keeping a DAS front‑end and adding a light U‑Net back‑end we can:
- Keep end‑to‑end with low latency for real time applications.
- Learn to sharpen beams and suppress artifacts.
- Output detection, localization, and class labels in one shot — matching the three functional blocks proposed by [1].
3 METHODOLOGY
3.1 System Overview
Figure 1 shows an overview of the proposed system pipeline. It consists of a custom microphone array for capturing multichannel audio, DAS beamforming to generate spatial energy maps, dataset construction with GPS-based labeling, a U-Net segmentation model, and a centroid-based post-processing step for estimating the direction of arrival (DOA).
3.2 Microphone Array Design and Recording Setup
We assembled a 24-channel microphone array using six Rode microphones mounted on each of the three legs of a standard tripod, forming an upright tetrahedral array as shown in Fig. 2.
The remaining six microphones were arranged in a horizontal circular ”yellow” ring. The array geometry provides progressive inter-microphone distances ranging from 4 cm to 1.1 m, optimizing for spatial aliasing and directional resolution in the frequency band between approximately 200 Hz and 4000 Hz, based on the array aperture and the speed of sound. Four Zoom F6 multichannel recorders (6 channels each) were synchronized via a synthetic impulse: channel 1 of every unit is placed at the array origin, the impulse is played once, and all files are shifted until their peak samples coincide, yielding ±1 sample at 48 kHz (7 mm acoustic error). An Insta360 X4 camera was mounted on the top to capture visual reference of drone takeoff and synchronize audio and flight logs temporally.
3.3 Field Recordings and Data Acquisition
A total of six open-field recording sessions were conducted using a DJI Air 3 drone, across three different dates and two distinct locations. Each session produced 24-channel synchronized audio recordings signals in .wav format, 360° video files, and GPS log files from the drone’s onboard telemetry. The GNSS (Global Navigation Satellite System) on the DJI Air 3 provides ±0.5 m position accuracy. An additional log file was recorded with the drone stationary at the array’s location to define the Cartesian origin for GPS transformation and reference position. This reference enabled accurate computation of azimuth and elevation angles relative to the array.
3.4 Beamforming Map Construction and Spectral Feature Extraction
Each GPS log point from the drone is transformed to a Cartesian coordinate system with the microphone array origin at . The azimuth () and elevation () are then calculated as:
For beamforming, the DAS output in a direction is computed by delaying the signals at each microphone by a steering delay and summing, N is the number of microphones:
In general, beamforming relies on the far-field assumption that the sources are far away and the waves become spherical or planar at the array position [2]. Given the position vector of each microphone in the array, the delay is obtained via the projection of onto the direction vector:
where is the speed of sound in air (typically chosen as 343 m/s). The signals are then shifted by these delays, and their average yields the beamformed output for a given steering direction. This approach is repeated for all directions in a discretized spatial grid to construct a full acoustic energy map. The acoustic beamformed maps were initially computed over a 2D rectangular grid covering azimuth angles in and elevation angles in , using a resolution of in both dimensions. For each azimuth-elevation pair, a DAS beamforming algorithm was applied to align and sum time-delayed microphone signals over a 100 ms window, sampled at 48 kHz. This yields a time-domain waveform of 4800 samples per spatial direction, producing a 3D tensor snapshot with shape , where is the number of azimuth bins, the number of elevation bins, and the number of time samples.
To convert these maps into a spectral representation, each time-domain waveform is transformed using the FFT. Only the 200–2200 Hz band is retained, corresponding to the dominant energy of the DJI Air 3 drone. The spectrum is uniformly divided into bins, which are globally normalized across spatial directions per frame. The result is a tensor of shape .
To better align with the spherical nature of the acoustic scene and reduce distortion in convolutional layers, the data is reprojected into a polar grid. In this transformation, elevation is mapped to radial distance from the center (with at the center and at the outer edge), and azimuth is mapped to angular position around the circle. The resulting spatial layout is a square grid of size , where the angular geometry is preserved. The final input tensor used for learning has shape .
3.5 Dataset Construction and GPS-Based Labeling
The dataset was recorded in multiple sessions. The training set comprises 30 minutes of drone flight and 10 minutes of ambient noise recorded in March 2025, along with an additional 22 minutes of drone flight from November 2024, all at Site 1. The test set consists of 20 minutes of drone flight and 10 minutes of ambient noise from March 2025, also at Site 1. The validation set includes 3 minutes of drone flight and 1 minute of ambient noise from October 2024, recorded at Site 2. Each flight session is treated as an independent data segment to prevent information leakage during training and evaluation.
Synchronized GPS logs from the drone are converted into spherical coordinates relative to the microphone array origin. Each 100 ms frame is annotated using a radial angular threshold around the ground-truth direction of arrival (DOA): all pixels within this threshold are labeled as 1 (drone-present), and the rest as 0 (drone-absent). This creates binary segmentation masks .
This labeling strategy compensates for beamforming limitations, such as reduced spatial resolution at low frequencies (due to wider beamwidth) and the presence of side lobes at high frequencies. By allowing spatial tolerance, the model is encouraged to learn smoother and physically grounded segmentation masks, consistent with prior studies on acoustic source mapping [27].
3.6 Dataset Representation
Each training example is a pair where the input tensor encodes the beamformed spectral information in polar coordinates, and the label is the corresponding binary segmentation mask.
3.7 U-Net Architecture and Hyperparameter Optimization
We implemented a modified U-Net architecture that accepts rectangular or polar input tensors with shape and outputs a binary segmentation mask . The architecture consists of an encoder with downsampling convolutional blocks, followed by a symmetric decoder with upsampling layers. Skip connections bridge corresponding encoder and decoder levels to preserve any spatial details lost during the downsampling. Optional attention gates [31] can be applied to skip connections, the bottleneck, or both. The number of base filters, depth of the encoder–decoder path, kernel size, and attention configuration are all tunable hyperparameters.
To optimize the model configuration, we used the Optuna framework to search the hyperparameter space. The best configuration was selected according to the minimum validation loss: We found 16 FFT bins, 64 base filters, depth 3, 3×3 kernels, lr = 0.005, and skip‑attention.
To address label sparsity and class imbalance, we selected the Tversky loss [32] which extends Dice loss by weighting false positives and is effective for small positive regions.
3.8 Inference and Evaluation
At test time, the U-Net outputs are thresholded, and a centroid is computed over active regions in the predicted mask to estimate the DOA. To suppress spurious activations we apply an erosion, retain the largest connected component and compute its centroid. Metrics are computed per 100 ms frame, then averaged over full trajectories to expose range-dependent accuracy.
4 RESULTS
The results were obtained using two test datasets: Test 2, collected in March 2025 at the original site, and Test 1 in October 2024, recorded at a different location and time under unseen conditions using the same DJI Air 3 drone and no-drone scenarios.
On the left side of the Fig. 3, the beamforming energy map is presented as a function of azimuth and elevation. The black X indicates the ground truth position of the drone (Log ref), extracted from synchronized GPS flight logs, while the red dot shows the location of maximum energy in the beamforming map (BF). It is visually evident that the beamforming-based estimation does not align with the true drone position.
On the right side, the U-Net creates binary segmentation output is displayed in polar coordinates. The black squares represent the segmented region predicted by the model. The blue triangle corresponds to the centroid of the predicted region. It can be observed that the U-Net model provides a significantly more accurate estimation of the drone’s position, closely matching the ground truth at a distance at 101 meters between the drone and the microphone array.
4.1 Performance Analysis by Distance Range
False Negative Rate (FNR). Fig. 4 shows the FNR as a function of source-array distance, grouped into three bins: 0–50 m, 50–100 m, and 100–200 m. A prediction is classified as a false negative if no output is generated or if the predicted centroid standard deviation exceeds over the last 3 observations, indicating high spatial uncertainty. In both datasets, the beamforming approach exhibits higher FNRs across all bins, particularly in the 100–200 m range, where signal energy is weaker. In contrast, the U-Net segmentation model maintains a lower FNR in many cases, suggesting greater robustness to acoustic attenuation and environmental variability.
Angular Error. Fig. 5 reports the mean angular error in degrees for each method and dataset. Beamforming shows increasing error with distance, especially beyond 100 m. The U-Net model consistently outperforms beamforming, with lower average angular errors across all distance bins. Notably, in the 50–100 m bin—where localization tends to be challenging—the model shows stable and accurate predictions even under environmental mismatch in the Test 1 dataset.
During a five‑minute recording with no drone present, we compared the false‑positive rates. Because there is no target signal in this scenario, performance was evaluated by measuring the standard deviation. Beamforming produced a 67.0% false‑positive rate, whereas the U‑Net reduced this to just 14.9%, demonstrating a substantial improvement.
Furthermore, both the false negative rate and the angular localization error can be mitigated by employing multiple synchronized devices with higher microphone density. This is possible within our framework due to the use of low-cost hardware components, enabling scalable deployments. These results confirm that the U-Net-based approach generalizes better than conventional beamforming across varying test conditions and distances.
5 Conclusion and Future Work
This work presents a U-net framework for sound source localization (SSL) that reinterprets beamformed acoustic maps as spatial segmentation tasks over the spherical field. By applying U-Net-based convolutional architectures to azimuth-elevation representations of DAS beamformed audio, we address classical limitations in beamforming—such as low-frequency blurring and high-frequency side lobes—thus enhancing angular resolution and robustness. Transforming beamformed maps into polar coordinates further aligns the spatial layout with spherical geometry, reducing distortion and better supporting CNN-based learning, as emphasized in DeepWave [10].
Unlike end-to-end models requiring raw microphone inputs or fixed array geometries, U-net model applies preprocessed spatial inputs, enabling different mic-array configurations. Combined with real-world drone recordings and GPS-based supervision, the proposed method demonstrates generalization across distances and locations, offering a practical solution for low-latency acoustic perception.
Future work will focus on extending the system to handle spatio-temporal dynamics via recurrent or attention-based models like Mark R-CNN[33] or yolact++[34] and VisTR[35], supporting multi-source and multiclass segmentation, and using synthetic data generation pipelines for diverse acoustic environments.
By framing SSL as a semantic segmentation task on beamformed maps, this approach bridges spatial signal processing and computer vision, opening new directions for high-resolution, real-time acoustic scene understanding.